OpenStudy (softballgirl372015):

Aglebra 2 Help!! If P(x)=x^3-4x^2+cx+6 , for what value of c is P(3)=0?

OpenStudy (softballgirl372015):

do you have any idea how to solve this?

OpenStudy (radar):

Plug in the 3 getting: 3^3 - 4 (3^2) + 3c + 6 = 0 \[3^{3} - 4(3^{2}) + 3c + 6=0\]

OpenStudy (radar):

Can you now solve for c ?

OpenStudy (softballgirl372015):

so c=1 ?

OpenStudy (radar):

No. working it out 27 - 36 + 3c + 6 = 0 combine and solve.

OpenStudy (softballgirl372015):

so what is the answer?

OpenStudy (softballgirl372015):

i thought that it would be 1

OpenStudy (radar):

combine 27-36+6 can you add signed numbers?

OpenStudy (softballgirl372015):

yes i can it equals -3

OpenStudy (radar):

LOL I goofed so -3 + 3c=0 and 3c=3 and c does indeed = 1 Sorry I was wrong.

OpenStudy (softballgirl372015):

its okay! can you please help me with another problem?

OpenStudy (radar):

Yes, if I can learn how to add signed numbers.

OpenStudy (softballgirl372015):

LOL thanks! IM going to type it right now.

OpenStudy (softballgirl372015):

If a number is selected at random from {-4, (2+2i)^2, (3-i)^2, (2i)^6, 0, (negative sqrt 3 -isqrt3)^2} what is the probability that the number is real?

OpenStudy (softballgirl372015):

the i is an imaginary number

OpenStudy (radar):

Wow that course is something else, they are testing you on probability and the ability to check out for real numbers from complex ones. First would you agree that they have given you a total of 6 numbers (including 0)

OpenStudy (softballgirl372015):

yup!

OpenStudy (softballgirl372015):

i dont know if this answer is right but I got 1/2. Is that what you think it is?

OpenStudy (radar):

OK, 4 is real 2i^6 is real 0 is real Now for the (2+2i)^2 =4 + 8i +4i^2 = 8i imaginary (3-i)^2 =9 - 6i -1 =8 - 6i a complex number with imaginary component \[(-\sqrt{3}-i \sqrt{3})^{2}\] =6i which is imaginary

OpenStudy (softballgirl372015):

okay so then would you set up the fraction of 3/6 which simplifies to 1/2?

OpenStudy (radar):

Looks like half are real and half are complex 50% probability wouldn't you say, in fact you did say. Do you have a choice for 50% or 1/2 or is this a multiple choice problem?

OpenStudy (softballgirl372015):

the 3 because there are 3 real roots and then 6 because there are 6 numbers that you can choose from.

OpenStudy (softballgirl372015):

it is an open ended question, but this makes sense to me.

OpenStudy (radar):

Well it looks like a good solution. What kind of school are you going to?

OpenStudy (softballgirl372015):

im a sophomore in high school taking honors algebra 2.

OpenStudy (softballgirl372015):

i dont want to use up all of your time but can you help me with 2 more problems? I solved one of them so far but am not sure if i have the correct answer.

OpenStudy (radar):

Very good, good luck in your studies, and in softball. My granddaughter got hurt in soft ball, but is all healed up now, she was playing second base, and a runner slid into second and hit her leg with cleats, took 50 stitches!

OpenStudy (radar):

OK, maybe I can

OpenStudy (softballgirl372015):

wow! 50 stitches I hope she recuperates well and does good this season!

OpenStudy (softballgirl372015):

the problem is to solve ix^2+5x+24i+0 for x. The i is imaginary numbers again.

OpenStudy (radar):

Is their an equal sign somewhere in there?

OpenStudy (softballgirl372015):

yes, ix^2+5x+24=0 SOrry about that. I solved the problem and got -3i and 8i but I am not sure if there are supposed to be two answers.

OpenStudy (softballgirl372015):

ix^2+5x+24i=0

OpenStudy (softballgirl372015):

sorry again, I wrote the wrong equation down

OpenStudy (radar):

Is the first term \[ix ^{2}........or (ix)^{2}\]

OpenStudy (softballgirl372015):

it is ix^2 where the x is raised to the second power

OpenStudy (radar):

That factors to (ix + 8) (x + 3i) is that how you worked it?

OpenStudy (radar):

ix + 8 = 0 ix = -8 multiply both sides by i -x = -8i x = 8i and the other root is x = -3i

OpenStudy (softballgirl372015):

Thats what I got as well! I used the quadratic formula though. Thank you sooo much for your help!!! You've really helped me a lot. Are you up for one more? LOL

OpenStudy (radar):

One more, and then I got to get off this computer lol

OpenStudy (softballgirl372015):

HAHAH!! Thank you!! Here is the problem: Let P(x)=ax^5+bx^3+cx, where a, b, and c are negative. Explain why P cannot have any negative zeros.

OpenStudy (softballgirl372015):

I tried to use Descartes Rule of Signs, but didnt know if I was going on the right path because when using that rule, I got that there would be negative zeros. So I am a little confused.

OpenStudy (radar):

I don't know why! Whats your idea. I can see that when x is positive that the terms would have all negative coefficients there fore negative values, If x is negative since they are all odd powers the value would be negative and with negative coefficients become positive and could result in negative roots ???????sorry about this, but I am also confused.

OpenStudy (radar):

I guess I am missing something here.

OpenStudy (softballgirl372015):

Thats what I was thinking. because when I used Descartes Rule of Signs, I resulted in 2 negative roots. Oh well. LOL. Thank you so much for everything that you have helped me with!! I feel confident about my math homework now! Thanks :)

OpenStudy (softballgirl372015):

I have one more question though....Just kidding! LOL :) Have a good night.

OpenStudy (radar):

You too, I don't mind trying to help when I see that the student is trying also. Again good luck.

OpenStudy (radar):

I realize you have closed this question, still I have been thinking about it. "Here is the problem: Let P(x)=ax^5+bx^3+cx, where a, b, and c are negative. Explain why P cannot have any negative zeros." Factoring getting P(x)= x(ax^4 + bx^2 +c) Note that one zero is x=0. Now the other factor has all even powers and with negative coefficients would all be negative resulting in positive roots. I think!?

OpenStudy (softballgirl372015):

Oh! I understand what you are saying! I've been trying to figure this out. When you factor the remaining quadratic, the roots are positive! I am going to plug it in right now. Thanks! You're awesome!

OpenStudy (softballgirl372015):

What you said makes sense to me, so I am trying to come up with an equation to test this.